OpenStudy (brooke10239):

Cody is selling chocolate and white chocolate candy bars for a school fundraiser. 30% of the candy bars sold have been white chocolate. If Cody has sold 24 white chocolate candy bars, how many total candy bars has he sold

OpenStudy (brooke10239):

A. 20 B. 40 C. 60 D. 80

OpenStudy (anthonygarrison875):

i think its c

OpenStudy (brooke10239):

Okay Thanks!

OpenStudy (quinnlover):

i like chocolate

OpenStudy (perl):

c) is wrong

OpenStudy (anthonygarrison875):

i doo to yum yum'

OpenStudy (brooke10239):

Oh, then what is it?

OpenStudy (quinnlover):

:D yassssss

OpenStudy (anthonygarrison875):

lol

OpenStudy (perl):

We can set up a proportion. 30/100 = 24/ x

OpenStudy (quinnlover):

it would be d right?

OpenStudy (perl):

$$\large \frac{30}{100} = \frac{24}{ x }$$